OpenStudy (anonymous):

the larger of two numbers is 6 less than five times the smaller. The sum of the two numbers is 102 what are the numbers?

OpenStudy (jlastino):

let the smaller number be x the larger number is 6 less than five times the smaller...how would you represent this in terms of x?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

5x-n

OpenStudy (anonymous):

Let n = your smaller number, set up an equation.

OpenStudy (anonymous):

the equation above is the one i wrote out an i'm nt getting 102 so i still dt kno how to set up the equation

OpenStudy (anonymous):

You have 2 variables in that equation, what do they represent?

OpenStudy (anonymous):

x is the larger number n is the smaller number that's all i know- but idk how to represent them hence the reason y i'm asking for help

OpenStudy (anonymous):

x+n = 102

OpenStudy (jlastino):

let the smaller number be x the larger number is 6 less than five times the smaller so the large number would 5x-6 the sum of these numbers equal to 102 x+5x-6=102 6x=108 x=18 5x-6=84 The two numbers are 18 and 84

OpenStudy (anonymous):

thanks-